meganmoney543:

Write the equation of a line, in general form, with a point (-3, 0) and an undefined slope?

meganmoney543:

x + 3 = 0 y + 3 = 0 x - 3 = 0 y - 3 = 0

daddysbabygirl:

which ones can you eliminate

meganmoney543:

idk

daddysbabygirl:

what's an undefined slope

meganmoney543:

i dont kno i slept thru this

daddysbabygirl:

An undefined slope (or an infinitely large slope) is the slope of a vertical line! The x-coordinate never changes no matter what the y-coordinate is! There is no run!

meganmoney543:

ok

daddysbabygirl:

Undefined slope is when the line is vertical, and the slope is infinite (technically); Any y value is irrelevant in this kind of line, so in slope intercept, the equation is x = -3.; Putting it in standard form, you get x + 3 = 0;

meganmoney543:

thxxxxxxx

daddysbabygirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @meganmoney543 thxxxxxxx \(\color{#0cbb34}{\text{End of Quote}}\) np